# Compact manifolds of dimension $n\geq 12$ with positive isotropic curvature

### v21

@@ 5 Appendix: gluing and strengthening of $\varepsilon$-necks
In particular, the metric $e^{-2\varphi}g_{\text{cly}}|_{\mathbb{S}^{n-1}\times(-\infty,B]}$ has constant sectional curvature $k^{2}$ on the region $\mathbb{S}^{n-1}\times[B^{\prime},B]$. Then we smoothly glue a metric $n$-ball of the same constant sectional curvature $k^{2}$ to $(\mathbb{S}^{n-1}\times(-\infty,B],e^{-2\varphi}g_{\text{cly}}|_{\mathbb{S}^{n-1}\times(-\infty,B]})$ along their boundaries, and the resulting Riemannian manifold is a possible choice for $(\mathbb{R}^{n},\hat{g}(0))$. Let $\Gamma$ be a finite subgroup of $O(n)$ acting freely on $\mathbb{S}^{n-1}$, and $(\mathbb{R}^{n}//\Gamma,\hat{g}_{\Gamma}(t))$ be the quotient of a smooth standard solution by $\Gamma$, called an orbifold standard solution. So the initial metric $\hat{g}_{\Gamma}(0)$ of an orbifold standard solution is obtained by gluing a (quotient) cap which is diffeomorphic to $D^{n}//\Gamma$ and also called the surgery cap, to the half-infinite quotient cylinder $\mathbb{S}^{n-1}/\Gamma\times(-\infty,0]$ with scalar curvature 1 along their boundaries. Let $p_{\Gamma}$ be the tip of an orbifold standard solution $(\mathbb{R}^{n}//\Gamma,\hat{g}_{\Gamma}(t))$, that is the image of the point $p_{0}$ under the natural projection $\pi_{\Gamma}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}//\Gamma$.
By the maximum principle $R(x,t)=R_{\hat{g}(t)}(x)\geq\frac{1}{1-\frac{2}{n}t}$ for any $x\in\mathbb{R}^{n}$. So if $t>\frac{n}{n+2}$, $tR(x,t)\geq t\frac{1}{1-\frac{2}{n}t}>1$. We’ll fix a $n$-dimensional smooth standard solution. Let $K_{\text{st}}$ be the superemum of the norm of the Riemannian curvature tensor of the $n$-dimensional smooth standard solution on the time interval $[0,\frac{n+1}{n+2}]$.
@@
Lemma 5.8.
(cf. Lemma 4.3.5 in [6]) Let $n\geq 3$, and $K=\max\{K_{\text{st}},1\}$. For any $\varepsilon\in(0,10^{-3})$ there exists a number $\beta=\beta(\varepsilon)(=\beta(\varepsilon,n))\in(0,10^{-2})$ with the following property. Let $a,b$ be real numbers with $a<b\leq 0$, $|b|\leq\frac{n}{n+1}$, and $(\mathcal{O}(\cdot),g(\cdot))$ be a $n$-dimensional complete surgical solution to the Ricci flow defined on $(a,0]$ with at most isolated singularities, and $x$ be a point in $\mathcal{O}(b)$ such that:
@@
Proof. The proof is adapted from that of Lemma 4.3.5 in [6]. We argue by contradiction. Otherwise there exist $\varepsilon\in(0,10^{-3})$, a sequence of positive numbers $\beta_{k}\rightarrow 0$, sequences $a_{k}<b_{k}$, $b_{k}\in[-\frac{n}{n+1},0]$, and a sequence of $n$-dimensional surgical solutions $(\mathcal{O}_{k}(t),g_{k}(t))$ ($t\in(a_{k},0]$) with a point $x_{k}\in\mathcal{O}_{k}(b_{k})$ such that
@@
Now we see that the sequence $(B_{\Psi_{k}^{*}g_{k}(0)}(*,c(\beta_{k}\varepsilon)^{-1}),\{\Psi_{k}^{*}g_{k}(t)\}_{t\in(b_{k}-Q_{k}^{-1},0]},(*,0))$ subconverges smoothly to $(\mathbb{S}^{n-1}\times\mathbb{R},\{\tilde{g}(t)\}_{t\in(b-1,0]},(*,0))$, where $\tilde{g}(t)=g_{\text{cly}}(t-b)$. Using assumption (ii), $Q_{k}\rightarrow 1$, and $b_{k}\rightarrow b$ we deduce that the tensor $\Psi_{k}^{*}g_{k}(b)-\tilde{g}(b)$ together with its the covariant derivatives are tending to 0 on arbitrarily large balls centered at $(*,b)$. Then by the argument in Section 2 of [46], we see that the diffeomorphisms involved in the subconvergence here can be chosen to be the identity maps. So if $k$ is sufficiently large, the evolving tensor $\Psi_{k}^{*}g_{k}(\cdot)-\tilde{g}(\cdot)$ together with all its covariant derivatives are arbitrarily close to 0 on arbitrary compact subsets of $\mathbb{S}^{n-1}\times\mathbb{R}\times(b-1,0]$ containing $(*,0)$. As $\Psi_{k}$ is $\Gamma_{k}$-invariant, it follows from Remark 4.2.7(i) in [6] and Lemma 5.3 that (a suitable restriction of) the diffeomorphism $\psi_{k}$ will give a strong $\varepsilon$-neck centered at $(x_{k},0)$ for sufficiently large $k$. This contradicts assumption (iv). $\Box$
Note that the assumption of our Lemma 5.8 is slightly weaker than that of Lemma 4.3.5 in [6]. That is, in assumption (ii) we do not assume that the $\beta\varepsilon$-neck is strong, which is assumed in Lemma 4.3.5(ii) in [6]. (Of course, we need to adjust assumption (iii) compared to Lemma 4.3.5(iii) in [6].)

### v22

@@ Appendix A Appendix: gluing and strengthening of $\varepsilon$-necks
In particular, the metric $e^{-2\varphi}g_{\text{cly}}|_{\mathbb{S}^{n-1}\times(-\infty,b_{0}]}$ has constant sectional curvature $k^{2}$ on the region $\mathbb{S}^{n-1}\times[b^{\prime},b_{0}]$. Then we smoothly glue a closed metric $n$-ball of the same constant sectional curvature $k^{2}$ to $(\mathbb{S}^{n-1}\times(-\infty,b_{0}],e^{-2\varphi}g_{\text{cly}}|_{\mathbb{S}^{n-1}\times(-\infty,b_{0}]})$ along their boundaries, and the resulting Riemannian manifold is a possible choice for $(\mathbb{R}^{n},\hat{g}(0))$. Let $\Gamma$ be a finite subgroup of $O(n)$ acting freely on $\mathbb{S}^{n-1}$, and $(\mathbb{R}^{n}//\Gamma,\hat{g}_{\Gamma}(t))$ be the quotient of a smooth standard solution by $\Gamma$, called an orbifold standard solution. So the initial metric $\hat{g}_{\Gamma}(0)$ of an orbifold standard solution is obtained by gluing a (quotient) cap which is diffeomorphic to $D^{n}//\Gamma$ and also called the surgery cap, to the half-infinite quotient cylinder $\mathbb{S}^{n-1}/\Gamma\times(-\infty,0]$ with scalar curvature 1 along their boundaries. Let $p_{\Gamma}$ be the tip of an orbifold standard solution $(\mathbb{R}^{n}//\Gamma,\hat{g}_{\Gamma}(t))$, that is the image of the point $p_{0}$ under the natural projection $\pi_{\Gamma}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}//\Gamma$.
By the maximum principle $R(x,t)=R_{\hat{g}(t)}(x)\geq\frac{1}{1-\frac{2}{n}t}$ for any $x\in\mathbb{R}^{n}$. So if $t>\frac{n}{n+2}$, $tR(x,t)\geq t\frac{1}{1-\frac{2}{n}t}>1$. We’ll fix an $n$-dimensional smooth standard solution. Let $K_{\text{st}}$ be the superemum of the norm of the Riemannian curvature tensor of the $n$-dimensional smooth standard solution on the time interval $[0,\frac{n+1}{n+2}]$.
@@
Lemma A.8.
(cf. Lemma 4.3.5 in [6]) Let $n\geq 3$, and $K=\max\{K_{\text{st}},1\}$. For any $\varepsilon\in(0,10^{-3})$ there exists a number $\beta=\beta(\varepsilon)(=\beta(\varepsilon,n))\in(0,10^{-2})$ with the following property. Let $a,b$ be real numbers with $a<b\leq 0$, $|b|\leq\frac{n}{n+1}$, and $(\mathcal{O}(\cdot),g(\cdot))$ be an $n$-dimensional complete surgical solution to the Ricci flow defined on $(a,0]$ with at most isolated singularities, and $x$ be a point in $\mathcal{O}(b)$ such that:
@@
Proof. The proof is adapted from that of Lemma 4.3.5 in [6]. We argue by contradiction. Otherwise there exists $\varepsilon\in(0,10^{-3})$, a sequence of positive numbers $\beta_{k}\rightarrow 0$, sequences $a_{k}<b_{k}$, $b_{k}\in[-\frac{n}{n+1},0]$, and a sequence of $n$-dimensional surgical solutions $(\mathcal{O}_{k}(t),g_{k}(t))$ ($t\in(a_{k},0]$) with a point $x_{k}\in\mathcal{O}_{k}(b_{k})$ such that
@@
Now we see that the sequence $(B_{\Psi_{k}^{*}g_{k}(0)}(*,c(\beta_{k}\varepsilon)^{-1}),\{\Psi_{k}^{*}g_{k}(t)\}_{t\in(b_{k}-Q_{k}^{-1},0]},(*,0))$ subconverges smoothly to $(\mathbb{S}^{n-1}\times\mathbb{R},\{\tilde{g}(t)\}_{t\in(b-1,0]},(*,0))$, where $\tilde{g}(t)=g_{\text{cly}}(t-b)$. Using assumption (ii), $Q_{k}\rightarrow 1$, and $b_{k}\rightarrow b$ we deduce that the tensor $\Psi_{k}^{*}g_{k}(b)-\tilde{g}(b)$ together with its the covariant derivatives are tending to 0 on arbitrarily large balls centered at $(*,b)$. Then by the argument in Section 2 of [49], we see that the diffeomorphisms involved in the subconvergence here can be chosen to be the identity maps. So if $k$ is sufficiently large, the evolving tensor $\Psi_{k}^{*}g_{k}(\cdot)-\tilde{g}(\cdot)$ together with all its covariant derivatives are arbitrarily close to 0 on arbitrary compact subsets of $\mathbb{S}^{n-1}\times\mathbb{R}\times(b-1,0]$ containing $(*,0)$. As $\Psi_{k}$ is $\Gamma_{k}$-invariant, it follows from Remark 4.2.7(i) in [6] and Lemma A.3 that (a suitable restriction of) the diffeomorphism $\psi_{k}$ will give a strong $\varepsilon$-neck centered at $(x_{k},0)$ for sufficiently large $k$. (By the way, Remark 4.2.7(i) in [6] follows from the fact that given $n>1$, $\varepsilon>0$ and $Q>0$, there exists $\hat{Q}>Q^{-1}$ such that
$\sup_{t\in[-1,0]}|\hat{Q}Qg_{\text{cyl}}((\hat{Q}Q)^{-1}t)-g_{\text{cyl}}(t)|_{[\varepsilon^{-1}],\mathbb{S}^{n-1}\times(-\frac{1}{\varepsilon},\frac{1}{\varepsilon}),g_{\text{cyl}}(t)}<\varepsilon,$
using the notation in Definition 2.2.4 in [6].) This contradicts assumption (iv). $\Box$
Note that the assumption of our Lemma A.8 is slightly weaker than that of Lemma 4.3.5 in [6]. That is, in assumption (ii) we do not assume that the $\beta\varepsilon$-neck is strong, which is assumed in Lemma 4.3.5(ii) in [6]. (Of course, we need to adjust assumption (iii) compared to Lemma 4.3.5(iii) in [6].)